Ships A and B leave port together. For the next two hours, ship A travels at 20.0mph in a direction 70.0 ? west of north while the ship B travels 35.0 ?east of north at 20.0mph .

What is the distance between the two ships two hours after they depart?

What is the speed of ship A as seen by ship B?
Answer #1

XA = 2 * 20 sin70 = 37.58 miles west
YA = 2 * 20 cos30 = 13.6 miles north

XB = 2 * 20 cos35 = 22.94 miles east
YB = 2 * 20 sin35 = 32.766 miles north

Distance difference = (60.52, -19.166) = 63.48 miles

The speed is just the distance divided by 2 hours: 31.74mph